OpenStudy (anonymous):

If g(x) = x^2 − 3, find g(4). My answer is 13 Am I right? :)

OpenStudy (anonymous):

Plug in 4 in place of x: g(x) = 4^2 - 3 = 16 - 3

OpenStudy (anonymous):

yr ans is correct

jimthompson5910 (jim_thompson5910):

yes you are g(x) = x^2 - 3 g(4) = 4^2 - 3 g(4) = 16 - 3 g(4) = 13

OpenStudy (anonymous):

Thanks you! :)